# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v1

@@ 3.1 Challenges
Table 2 provides a detailed summary of the detection and diagnosis challenges specific to each competition. While each competition faces unique challenges due to its distinct industry background and objectives, a common issue they all share is the need to accurately classify multiple fault modes. Apart from the common issue, notably, the PHM EU competition in 2021 and the PHM AP in 2023 are both constrained by their small datasets, with the former comprising only 70 training and 29 test samples, and the latter having 177 training and 46 test samples. This paucity of data presents substantial obstacles for training machine learning models due to limited information from data. Additionally, the class imbalance present in the datasets of the 2021 and 2022 PHM EU competitions shows a significant imbalance with a higher number of healthy experiments compared to the faulty ones, thus impeding the models’ ability to detect rare fault conditions and challenging their robustness and generalization potential. In the meantime, these datasets also faced the obstacle of missing data, necessitating imputation techniques for robust data handling. The 2022 PHM NA competition confronts a domain-shift problem, where varying operational conditions and data from different assets result in inconsistent signal distributions, which complicates the task of accurate fault classification. Lastly, the 2023 PHM AP competition includes unexpected fault modes in the testing data, absent from the training set, reflecting the unpredictability of real-world industrial settings and underscoring the necessity for constructing robust and generalized models.
@@
Table 2: Overview of Detection and Diagnosis Challenges in PHM Data Competitions
2021 PHM EU 2022 PHM EU 2022 PHM NA 2023 PHM AP 2023 IEEE
Failure Mode 8 1+1+2 10 3 4
Asset NA NA 8 4 NA
Operating Condition 2 1+1+1 3 1 2
Volume of Data Limited Large Large Limited Large
(70 Train, 29 Test) (SPI:5985382, AOI:31617) (37229 Train, 16396 Test) (177 Train, 46 Test) (Total 50000)
Sampling Rate 0.1 Hz NA 50 kHz 1 kHz 10 kHz
Sampling Interval 1-3 hours NA NA 1.2 s 5 minutes
Table 3: Overview of Detection and Diagnosis Problems in PHM Data Competitions
@@
In a collaborative effort with the Swiss Centre for Electronics and Microtechnology (CSEM), a data challenge competition was introduced, offering a dataset derived from a real-world industrial manufacturing line dedicated to testing electrical fuses. The objective of this competition is to perform fault identification and classification, root cause analysis, rapid prediction, and system operation parameter identification. The dataset showed eight unique system failure modes, replicated under two distinct operating conditions. However, its scale was rather restricted, comprising merely 70 training and 29 testing samples. The dominant challenge with the training data was its class imbalance—most samples depicted a healthy condition. This scenario, coupled with the limited volume of data, presented challenges to the data-driven approach.
@@
Amidst this backdrop, the winning solution was a notable combination of decision tree algorithms and a propagation system [12]. While the decision trees zeroed in on observational diagnoses, the propagation system skillfully tackled chronological intricacies by infusing a Kalman-style filter. This fusion bolstered the reliability of the outcomes. To address the prominent data imbalance, the SMOTE (Synthetic Minority Oversampling Technique) method was utilized. In a related exploration, Ince et al. [24] incorporated the Leave One Feature Out Importance (LOFO-Importance) package for discerning essential features. Subsequently, they applied linear discriminant analysis for dimensionality reduction. During the modeling phase, different ML methods were employed — gradient boosting algorithms such as XGBoost and LightGBM, LDA classifier, and Gaussian process classifier. Especially for gradient boosting, they used Genetic Algorithms (GA) to optimize hyperparameters. A pivotal observation was XGBoost’s superior performance over the other algorithms. Diverging in their approach, Aimiyekagbon et al. [2] showcased a rule-based diagnostic technique, comparing its performance with that of Decision Trees and Random Forest methods. On the other hand, Aydemir et al. [4] proposed a regularized LSTM for sifting through vital features and then leveraged an ensemble of binary LSTM classifiers for fault detection and classification.
@@
Moreover, the tree-based algorithm XGBoost found favor not just in the aforementioned studies but also in three other distinct research papers [3, 55, 71] outside the competition. The distinctions among these studies were as follows: For instance, Tian et al. [71] relied on a feature importance ranking (FIR) method, targeting enhanced performance and simplification in intricate industrial classification scenarios, whereas Ramezani et al. [55] zeroed in on the challenge of missing value imputation, harnessing Partial Least Squares (PLS-MV). A significant contribution of Ramezani’s framework was its explainability, achieved by pinpointing key performance indicators for each fault family using the SHAP method [48]. In conclusion, tree-based methods can significantly enhance fault diagnosis, thereby boosting the efficiency and safety of manufacturing production lines.
@@
The 2022 European PHME Data Challenge, hosted in collaboration with Bitron Spa, spotlights a classification issue within an actual industrial Printed Circuit Board (PCB) production line. The challenge’s objectives span three pivotal sub-tasks: predicting Automatic Optical Inspection (AOI) defect detection based on Solder Paste Inspection (SPI) data, predicting human-made visual inspection labels (OperatorLabel), and finally, predicting the human-assigned repair label (RepairLabel). It’s noteworthy that, among the nine datasets discussed in this paper, this dataset stands unique. While others are time series data, this one is parametric data. Due to the extensive dataset, tree-based and neural network-based machine learning methods usually come to the fore.
@@
Delving into the competition’s outcomes, Gaffet’s team clinched 1st place with an XGBoost method based on encoding and feature engineering [18]. Notably, they further harnessed the SHAP method [48] for model interpretation. Echoing a similar strategy, Taco et al. [68] leveraged two tree-based gradient boosting algorithms, namely LightGBM and XGBoost. Their approach centered on solving classification problems by extracting significant statistical data during the feature engineering phase. The importance of feature engineering was further emphasized by Tang et al. [70]. They introduced a novel statistical feature extraction method coupled with a PinNumber-based technique. This method aimed to compress pin-level data into component-level information. When predicting automatic inspection defects, they integrated a neural network model, factoring in feeding imbalance control to navigate data imbalance challenges. Their toolkit also boasted a random forest model for both human inspection and repair predictions. Schmidt’s team [59], though paralleling Tang et al. in several aspects, differentiated their approach in ML model construction. Apart from their extensive focus on feature engineering, they applied a multi-layer perceptron neural network for defect predictions in automated inspections. For human inspection outcomes, a random forest algorithm was their choice, while decision trees were favored for predicting human repairs.
@@
Outside of competition, Mirzaei’s team delved deep into the challenge of imbalanced data [51]. Their solution entailed a data-level technique that integrated Recursive Feature Elimination (RFE) for feature selection and oversampling methods. This blend ensured balanced representation for minority classes, enhancing the performance of multiple ML algorithms, from Decision Trees and Random Forests to SVMs and 1dCNNs. In a somewhat divergent approach, Lee et al. [37] unveiled a new quality management paradigm termed Stream-of-Quality (SoQ) tailored for multi-stage manufacturing processes. By leveraging this dataset, they showcased the potency of their methodology. The implications of their work are vast, offering promising avenues for refining industrial AI algorithms and methodologies in a systematic manner.
@@
Rock drills play a pivotal role in sectors like mining, tunneling, and construction. Due to the potential economic and human costs from work interruptions caused by rock drill faults, ensuring accurate fault diagnosis is imperative. This data challenge tackles the intricate task of fault classification within various rock drill configurations. Challenges arise due to the domain-shift phenomenon, stemming from data sourced from diverse rock drill machines, which results in heterogeneous signal distributions that potentially impair classification accuracy. The dataset [27] encompasses ten distinct fault modes and one health mode, with training, validation, and testing data sizes of 34,045, 3,184, and 16,396, respectively.
@@
In the competition, the top accolade was clinched by Oh et al. [53], who championed a hybrid strategy, melding data-driven techniques with signal-processing methods. Their approach harnessed domain adaptation, metric learning, and pseudo-label-based deep learning to construct an ensemble deep learning model for comprehensive fault classification. For samples that posed challenges for deep learning, they deployed signal processing methods like Dynamic Time Warping (DTW), and Cross-correlation, and used SVM for supervised learning. The runner-up solution introduced a data-cropping technique, employing a convolutional neural network (CNN) as a pivotal feature extractor to bridge data length discrepancies. Kim et al. [31] innovatively addressed the domain-shift issue through a domain-adaptation-based scheme that harnessed a domain adversarial learning neural network for extracting domain invariant features while using maximum mean discrepancy (MMD) minimization for bridging distribution discrepancy, and a soft voting ensemble to reduce model uncertainty. Moreover, Minami and team proposed an ensemble approach, appending specialized models onto a baseline (or initial) model. This ensemble incorporated domain adaptation strategies to accommodate domain fluctuations. The baseline model employed conventional ML algorithms like SVM, Random Forest, and XGBoost for whole multi-class classification, whereas the specialized sub-models, leveraging CNN for feature extraction and classification, concentrated on binary classifications targeting specific classes that poorly performed on the baseline model [50].
@@
Outside the competition, two notable data-driven strategies emerged. Ling’s team [43] deployed an end-to-end fault classification framework derived from X-Vectors [63], achieving integrated optimization of both feature extraction and classification phases. Taco et al. [69] proposed a novel deep forest algorithm that fused multi-grained scanning for feature extraction and a cascade forest for layered predictions to perform failure model classification. In summary, all aforementioned methods emphasize the importance of tackling the domain shift challenge. Utilizing domain adaptation and machine learning techniques can enhance the accuracy of fault detection and diagnosis in rock drills, ensuring safer and more reliable operations.
@@
Table 4: Overview of Detection and Diagnosis Methodologies in PHM Data Competitions
@@
In the competition, the champions Takanobu Minami and Jay Lee [49], introduced a novel two-step approach. Initially, a similarity measures-based model was proposed for the categorization of data into four distinct states. Subsequently, for data corresponding to the solenoid valve fault, a model incorporating physics-inspired features was employed to pinpoint the fault location and estimate the valve opening ratio. A noteworthy aspect of their methodology entailed the employment of Dynamic Time Warping (DTW) [7] on the training dataset. This step was instrumental in quantifying the variability across various segments of the sensor data. Based on data analysis, authors opted to leverage solely the initial valve opening data segment for feature extraction, thereby mitigating variations arising from individual or sample variations. This strategic refinement substantially improved the data quality, resulting in better model performance in fault detection and classification, with the team achieving 100% classification accuracy in the final competition. Standing in the second position, Lee et al. [40] devised a hybrid approach combining the XGBoost-based method and the rule-based method. While the XGBoost-based approach primarily addressed comprehensive fault classification, the rule-based method was employed to formulate the solenoid valve opening ratio equation. This was accomplished through polynomial fitting, rooted in intrinsic physical characteristics, enabling a precise estimation of the solenoid valve opening ratio. Hot on their heels, Kato et al. [29] utilized the K-NN algorithms to classify faults and pinpoint the location of anomalies. Prior to the classification, a differentiation between normal and anomalous data was executed based on a similarity method. Their estimation metrics for the valve opening ratio were hinged upon the similarity of time series waveforms. Moreover, Aimiyekagbon et al. [1] employed an ensemble framework, integrating K-means clustering and decision trees. This model, enriched by domain-specific expertise, exhibited commendable precision in both anomaly detection and fault diagnosis. A synthesis of the competition’s myriad approaches underscores the pivotal role of similarity-based methods and the extraction of physical features. These elements stand out as potent tools in the realm of fault detection and classification, especially when contending with limited datasets.
@@
Multiple fault classification under different operating conditions using vibration signals. Gearbox malfunctions in industrial machinery can result in significant downtime and damage, leading to substantial financial losses. Early detection and accurate diagnosis of these faults are paramount for ensuring the machinery’s safe operation, averting catastrophic failures, and minimizing maintenance costs. The objective of this competition [5] is to develop ML-based models that can efficiently detect faults in the planetary gearboxes of industrial machinery using vibration signals. The dataset [16] spotlights four prevalent sun gear faults: surface wear, chipped teeth, cracks, and missing teeth. Vibration signals have been recorded for a duration of five minutes each, with a sampling rate of 10 kHz, under two distinct operational conditions. Given the expansive data size, totaling 50,000 entries, the potential for deploying deep learning techniques is evident.
Various data-driven methodologies have been employed in this competition. For instance, Lee et al. [41] utilized ensemble-based CNN techniques to address fault detection via time-series vibration data. Their findings underscored the robust performance of integrating three convolution kernel-based methods such as ROCKET (RandOm Convolutional KErnel Transform) method [13], one-dimensional convolutional neural networks integrated with ResNet, and a fusion of LSTM and FCN in delivering classification results for multivariate time-series data. Similarly, both Shen et al. [60] and Matthias Kreuzer alongside Walter Kellermann [33], designed models based on residual-based CNNs techniques to navigate the fault classification challenge. Shen [60] incorporated the Short-Time Fourier Transform (STFT) [23], transforming frequency domain signals into time-frequency domain signals using Fourier analysis - an invaluable tool for interpreting time-evolving signals. On the other hand, Kreuzer and Kellermann [33] leveraged data augmentation and regularization techniques, enabling model construction with fewer parameters without sacrificing performance quality. In another noteworthy contribution, Sobha et al. [64] harnessed a tree classifier to select valuable features from raw vibration signals, subsequently developing a sequential neural network model tailored for the concurrent detection of multiple gear faults.

### v2

@@ 3.1 Challenges
Table 3: Overview of Detection and Diagnosis Challenges in PHM Data Challenge Competitions
@@
Table 3 provides a detailed summary of the detection and diagnosis challenges specific to each competition. While each competition faces unique challenges due to its distinct industry background and objectives, a common issue they all share is the need to accurately classify multiple fault modes. Apart from the common issue, notably, the PHM EU competition in 2021 and the PHM AP in 2023 are both constrained by their small datasets, with the former comprising only 70 training and 29 test samples, and the latter having 177 training and 46 test samples. This paucity of data presents substantial obstacles for training ML models due to limited information from data. Additionally, in the 2021 and 2022 PHM EU competitions, the datasets show a significant imbalance with a higher number of healthy experiments compared to the faulty ones, thus impeding the models’ ability to detect rare fault conditions and challenging their robustness and generalization potential to some degree. In the meantime, these datasets also faced the obstacle of missing data, necessitating imputation techniques for robust data handling. The 2022 PHM NA competition confronts a domain-shift problem, where varying operational conditions and data from different assets result in inconsistent signal distributions, which complicates the task of accurate fault classification. Lastly, the 2023 PHM AP competition includes unexpected fault modes in the testing data, absent from the training set, reflecting the unpredictability of real-world industrial settings and underscoring the necessity for constructing robust and generalized models.
@@
In a collaborative effort with the Swiss Centre for Electronics and Microtechnology (CSEM), a data challenge competition was introduced, offering a dataset derived from a real-world industrial manufacturing line dedicated to testing electrical fuses. The objective of this competition is to perform fault identification and classification, root cause analysis, rapid prediction, and system operation parameter identification. The dataset showed eight unique system failure modes, replicated under two distinct operating conditions. However, its scale was rather restricted, comprising merely 70 training and 29 testing samples. The dominant challenge with the training data was its class imbalance—most samples were in healthy condition. This scenario, coupled with the limited volume of data, presented challenges to the data-driven approach.
@@
Against this background, the winning solution was a combination of decision tree algorithms and a propagation system [22]. While the decision trees focused on observational diagnoses, the propagation system skillfully tackled chronological intricacies by infusing a Kalman-style filter. This fusion bolstered the reliability of the outcomes. To address data imbalance, the SMOTE (Synthetic Minority Oversampling Technique) method was utilized. Additionally, [41] incorporated the Leave One Feature Out Importance (LOFO-Importance) package for capturing essential features. Subsequently, they applied linear discriminant analysis (LDA) for dimensionality reduction. During the modeling phase, different ML methods were employed — gradient boosting algorithms such as Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LightGBM), LDA classifier, and Gaussian process classifier. Especially for gradient boosting, they used Genetic Algorithms to optimize hyperparameters. A pivotal observation was XGBoost’s superior performance over the other algorithms. Differing from their approach, [3] showcased a rule-based diagnostic technique, comparing its performance with that of decision trees and random forest methods. Additionally, [7] proposed a regularized LSTM for sifting through vital features and then leveraged an ensemble of binary LSTM classifiers for fault detection and classification.
@@
Moreover, the tree-based algorithm XGBoost found favor not just in the aforementioned studies but also in three other distinct research papers [4, 85, 106] outside the competition. The distinctions among these studies were as follows: For instance, [106] relied on a feature importance ranking (FIR) method, targeting enhanced performance and simplification in intricate industrial classification scenarios, whereas [85] focused on the challenge of missing value imputation, harnessing Partial Least Squares (PLS-MV). A significant contribution of Ramezani’s framework was its explainability, achieved by pinpointing key performance indicators for each fault family using the SHAP method [70]. In conclusion, tree-based methods can significantly enhance fault diagnosis, thereby boosting the efficiency and safety of manufacturing production lines.
@@
The 2022 European PHME Data Challenge, hosted in collaboration with Bitron Spa, spotlights a classification issue within an actual industrial Printed Circuit Board (PCB) production line. The challenge’s objectives span three pivotal sub-tasks: predicting Automatic Optical Inspection (AOI) defect detection based on Solder Paste Inspection (SPI) data, predicting human-made visual inspection labels (OperatorLabel), and finally, predicting the human-assigned repair label (RepairLabel). Task 1 and Task 2 are binary classification problems, whereas Task 3 is a multi-class classification problem in this data challenge. This explains why, in Table 2, the ”Failure Mode” is listed as ”1+1+2”. In Table 2, the ”Operating Condition” is listed as ”1+1+1”, which means each task has its own operating condition. It’s noteworthy that, among the nine datasets discussed in this paper, this dataset stands unique. While others are time series data, this one is parametric data.
@@
Delving into the competition’s outcomes, Gaffet’s team clinched 1st place with an XGBoost method based on encoding and feature engineering [32]. Notably, they further harnessed the SHAP method for model interpretation. Similarly, [102] leveraged two tree-based gradient boosting algorithms, namely LightGBM and XGBoost. Their approach centered on solving classification problems by extracting significant statistical data during the feature engineering phase. The importance of feature engineering was further emphasized by [104]. They introduced a novel statistical feature extraction method coupled with a PinNumber-based technique. This method aimed to compress pin-level data into component-level information. When predicting automatic inspection defects, they integrated a neural network model, factoring in feeding imbalance control to navigate data imbalance challenges. Additionally, a random forest model was developed for both human inspection and repair predictions. Similarly, apart from an extensive focus on feature engineering, [89] applied a multi-layer perceptron (MLP) neural network for defect predictions in automated inspections. For human inspection outcomes, a random forest algorithm was their choice, while decision trees were favored for predicting human repair labels.
@@
Outside of competition, Mirzaei’s team delved deep into the challenge of imbalanced data [75]. Their solution entailed a data-level technique that integrated recursive feature elimination (RFE) for feature selection and oversampling methods. This blend ensured balanced representation for minority classes, enhancing the performance of multiple ML algorithms, from decision trees and random forests to SVMs and 1dCNNs. In a somewhat divergent approach, [58] introduced a new quality management paradigm termed Stream-of-Quality (SoQ) tailored for multi-stage manufacturing processes. By leveraging this dataset, they showcased the effectiveness of their methodology. The implications of their work are vast, offering promising avenues for refining industrial AI algorithms and methodologies in a systematic manner.
@@
Rock drills play a pivotal role in sectors like mining, tunneling, and construction. Due to the potential economic and human costs from work interruptions caused by rock drill faults, ensuring accurate fault diagnosis is imperative. This data challenge tackles the task of fault classification within various rock drill configurations. Challenges arise due to the domain-shift phenomenon, stemming from data sourced from diverse rock drill machines, which results in heterogeneous signal distributions that potentially impair classification accuracy. The dataset [44] encompasses ten distinct fault modes and one health mode, with training, validation, and testing data sizes of 34,045, 3,184, and 16,396, respectively.
@@
In the competition, the champion was clinched by [80], who deployed a hybrid strategy, combining data-driven techniques with various signal-processing methods. Their approach harnessed domain adaptation, metric learning, and pseudo-label-based deep learning to construct an ensemble DL model for comprehensive fault classification. For samples that posed challenges for DL, they deployed signal processing methods like Dynamic Time Warping (DTW), and Cross-correlation, and used SVM for supervised learning. The runner-up solution introduced a data-cropping technique, employing a convolutional neural network (CNN) as a pivotal feature extractor to bridge data length discrepancies. [49] innovatively addressed the domain-shift issue through a domain-adaptation-based scheme that harnessed a domain adversarial learning neural network for extracting domain invariant features while using maximum mean discrepancy (MMD) minimization for bridging distribution discrepancy, and a soft voting ensemble to reduce model uncertainty. Moreover, Minami’s team proposed an ensemble approach, appending specialized models onto a baseline model. This ensemble incorporated domain adaptation strategies to accommodate domain fluctuations. The baseline model employed conventional ML algorithms like SVM, Random Forest, and XGBoost for whole multi-class classification, whereas the specialized sub-models, leveraging CNN for feature extraction and classification, concentrated on binary classifications targeting specific classes that poorly performed on the baseline model [74].
@@
Outside the competition, two notable data-driven strategies emerged.  [65] deployed an end-to-end fault classification framework derived from X-Vectors [95], achieving integrated optimization of both feature extraction and classification phases [65][103] proposed a novel deep forest algorithm that fused multi-grained scanning for feature extraction and a cascade forest for layered predictions to perform failure model classification. In summary, all aforementioned methods emphasize the importance of tackling the domain shift challenge. Utilizing domain adaptation and ML techniques can enhance the accuracy of fault detection and diagnosis in rock drills, ensuring safer and more reliable operations.
Table 4: Overview of Detection and Diagnosis Methodologies in PHM Data Challenge Competitions
@@
In the competition, the champions [73], introduced a novel two-step approach. Initially, a similarity measures-based model was proposed for the categorization of data into four distinct states. Subsequently, for data corresponding to the solenoid valve fault, a model incorporating physics-inspired features was employed to pinpoint the fault location and estimate the valve opening ratio. A noteworthy aspect of their methodology entailed the employment of DTW [11] on the training dataset. This step was instrumental in quantifying the variability across various segments of the sensor data. Based on data analysis, authors opted to leverage solely the initial valve opening data segment for feature extraction, thereby mitigating variations arising from individual or sample variations. This strategic refinement substantially improved the data quality, resulting in better model performance in fault detection and classification, with the team achieving 100% classification accuracy in the final competition. Standing in the second position, [61] devised a hybrid approach combining the XGBoost-based method and the rule-based method. While the XGBoost-based approach primarily addressed comprehensive fault classification, the rule-based method was employed to formulate the solenoid valve opening ratio equation. This was accomplished through polynomial fitting, rooted in intrinsic physical characteristics, enabling a precise estimation of the solenoid valve opening ratio. Additionally, [47] utilized the K-NN algorithms to classify faults and pinpoint the location of anomalies. Prior to the classification, a differentiation between normal and anomalous data was executed based on a similarity-based approach. Their estimation metrics for valve opening ratio were hinged upon the similarity of time series waveforms. Moreover, [2] employed an ensemble framework, integrating K-means clustering and decision trees. This model, enriched by domain-specific expertise, exhibited good precision in both anomaly detection and fault diagnosis. Various approaches in the competition underscore the pivotal role of similarity-based methods and the extraction of physical features. These approaches stand out as potent tools in the realm of fault detection and classification, especially when contending with limited datasets.
@@
Gearbox malfunctions in industrial machinery can result in significant downtime and damage, leading to substantial financial losses. Early detection and accurate diagnosis of these faults are paramount for ensuring the machinery’s safe operation, averting catastrophic failures, and minimizing maintenance costs. The objective of this competition is to develop ML-based models that can efficiently detect faults in the planetary gearboxes of industrial machinery using vibration signals. The dataset [29] spotlights four prevalent sun gear faults: surface wear, chipped teeth, cracks, and missing teeth. Vibration signals have been recorded for a duration of five minutes each, with a sampling rate of 10 kHz, under two distinct operational conditions. Given the expansive data size, totaling 50,000 entries, the potential for deploying DL techniques is evident.
Various data-driven methodologies have been employed in this competition. For instance, [62] utilized ensemble-based CNN techniques to address fault detection via time-series vibration data. Their findings underscored the robust performance of integrating three convolution kernel-based methods such as ROCKET (RandOm Convolutional KErnel Transform) method [23], one-dimensional convolutional neural networks integrated with ResNet, and a fusion of LSTM and Fully Convolutional Network (FCN) in delivering classification results for multivariate time-series data. Similarly, both [91] and [51], designed residual-based CNN models to address fault classification challenges. On the one hand, [91] incorporated the Short-Time Fourier Transform (STFT) [40], transforming frequency domain signals into time-frequency domain signals using Fourier analysis - an invaluable tool for interpreting time-evolving signals. On the other hand, [51] leveraged data augmentation and regularization techniques, enabling model construction with fewer parameters without sacrificing performance quality. In another noteworthy contribution, [96] harnessed a tree classifier to select valuable features from raw vibration signals, subsequently developing a sequential neural network model tailored for the concurrent detection of multiple gear faults.